OpenStudy (anonymous):

A pool that has 24,375 gallons of water is drained at the end of the summer. The water is drained at a rate of 75 gallons per minute. Let m = minutes. Let g = gallons of water in the pool. a) Write a linear equation that can be used to find the number of gallons of water in the pool after any number of minutes into the draining process. b) Use your equation to determine how much water will be in the pool after 20 minutes and after 1 hour. Explain how you arrived at your answer. c) Use your equation to determine how many hours it will take to empty the pool. Show your work and expla

OpenStudy (anonymous):

answer people!

OpenStudy (anonymous):

x = 24,375 - 75m When the pool is empy, x = 0 Your answer will be in minutes, for the last part of the problem, so you'll have to convert it to hours.

OpenStudy (anonymous):

can you figure out b and c for me :0

OpenStudy (anonymous):

Plug 20 minutes, then 60 minutes, into the equation. x is the amount of water left in the pool.

OpenStudy (anonymous):

is that for a,b, or c

OpenStudy (anonymous):

I explained C in my first response. Set x = 0, (no water left in the pool) and solve for m (the number of minutes that it takes)

OpenStudy (anonymous):

;) thnx

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

i am still confused lol

OpenStudy (anonymous):

sorry :(

OpenStudy (anonymous):

For part b, put 20 (for 20 minutes) into the equation for m: x = 24,375 - 75 *20 Your answer is the amount of water in the pool after 20 minutes. Then for one hour, put 60 minutes into the equation for m: x = 24,375 - 75*60

OpenStudy (anonymous):

and what is A?

OpenStudy (anonymous):

x = 24,375 - 75m

OpenStudy (anonymous):

so it will take 1 hour and 15 minutes to empty the pool?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

0 = 24,375 - 75m 75m = 24,375 m = 24,375/75 = 325 minutes = 5.4 hours